OpenStudy (anonymous):

If h(x) = ln(x) then h^-1(x) = ?

OpenStudy (anonymous):

e^x

OpenStudy (mertsj):

\[y=\ln x\] That is the same as: \[e^y=x\] And the inverse is : \[e^x=y\]

OpenStudy (mertsj):

So we see that the inverse of y = lnx is y = e^x